OpenStudy (anonymous):

hi what is the factor of 2x^2-3x-9?

OpenStudy (lgbasallote):

2x^2 - 3x - 9 break down the middle term 2x^2 + 6x - 3x - 9 group them (2x^2 + 6x) - (3x + 9) factor out 2x(x + 3) - 3(x+3) do you see it now?

OpenStudy (anonymous):

\[x^2-3x-9=0\] \[x^2-6x+3x-9=0\] \[2x(x-3)+3(x-3)=0\] now you try to simplify.

OpenStudy (anonymous):

you can use the quadratic formula i.e \[-b \pm \sqrt{b ^{2}-4ac}/2a\] here a=2 b=-3 c=-9 put the values of a, b and c in equation and you'll get your answer

OpenStudy (anonymous):

thanks for helping me

OpenStudy (anonymous):

factor of 2x+8

OpenStudy (anonymous):

28X+7

OpenStudy (anonymous):

we have to make the first term of these sums.

OpenStudy (anonymous):

I CANNOT SOLVE THAT ONE

OpenStudy (anonymous):

wait i am doing.

OpenStudy (anonymous):

hi what is the factor of 8x^3-216

OpenStudy (anonymous):

2x+342

OpenStudy (anonymous):

\[First Term= (Middle Term)^2/4*(Last Term)\]

OpenStudy (anonymous):

thanks,but im not asking what is the procedure im asking is what is the solution?

OpenStudy (anonymous):

kantutan

OpenStudy (anonymous):

gago ka talaga

OpenStudy (anonymous):

oo nga pala tuloy ung sa e nglish

OpenStudy (anonymous):

tar a abano Joke :D

OpenStudy (anonymous):

factor of 8x^3-216 \[8x(x^2-27)=0\] \[8x=0 \] \[x=0\] or \[x^2-27=0\] \[x^2=27\] \[x=\sqrt{27}\]

OpenStudy (anonymous):

abano add mo ko sa fwend m!!

OpenStudy (anonymous):

(x-6)(x+3)

OpenStudy (anonymous):

@danny_dumond Please don't post only answers as that's against the Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

well find something like (cx+a) (dx+b) to = 2x^2-3x-9

OpenStudy (anonymous):

Gosh sorry! :P